OpenStudy (nuccioreggie):

Jake has a bag of 50 beads, of which some are red and the remaining are yellow. Jake randomly pulls out a bead from the bag, records the color, and replaces it in the bag. Jake has already recorded 6 red and 9 yellow beads. Based on the result, what is most likely the number of red beads in the bag? 10 20 30 40

OpenStudy (geekfromthefutur):

Do you thin you know what it is?

OpenStudy (welshfella):

Hint the proportion of red beads pick was 6/(9+6) = 6/15 = 2/5

OpenStudy (nuccioreggie):

B @welshfella

OpenStudy (anonymous):

20

OpenStudy (anonymous):

6/15=2/5

OpenStudy (anonymous):

get it?